OpenStudy (anonymous):

classify each triangle as acute,right,or obtuse. the numbers are 3,8,10

OpenStudy (shadowfiend):

Do you know what each type of triangle means?

OpenStudy (anonymous):

yes basically the size of the angles

OpenStudy (shadowfiend):

Right. So if one of the angles is 90 degrees, it is a right triangle. If one of the angles is greater than 90 degrees, it is an obtuse triangle. Otherwise, it is an acute triangle (because all three angles are less than 90 degrees).

OpenStudy (anonymous):

wat? im supposed to see what 3,8,10 make as a triangle its geometry

OpenStudy (shadowfiend):

Ok, so the side lengths are 3 8 and 10.

OpenStudy (shadowfiend):

The easiest thing to check first is this: right triangles have sides that are related by the pythagorean theorem. So \(a^2 + b^2 = c^2\). \(c\) is always the hypotenuse, which is the longest side. Do these three numbers work in that equation?

OpenStudy (anonymous):

yea but it does not have a triangle its just ask question 7. 3,8,10 and its asking me the lenghts of the sides of a triangle are given.Classify each triangle as acute,right,or obtuse.

OpenStudy (shadowfiend):

I understand that. The first step is to check whether it's a right triangle by checking whether those side lengths work in the pythagorean theorem. Since 10 is the longest side, it would have to be the hypotenuse. So is it true that \(3^2 + 8^2 = 10^2\)?

OpenStudy (anonymous):

yes

OpenStudy (shadowfiend):

\(3^2 = 9\), \(8^2 = 64\), and \(10^2 = 100\). So 9 + 64 = 100?

OpenStudy (anonymous):

ok yes

OpenStudy (shadowfiend):

No, 9 + 64 is 73, which is not equal to 100. So this isn't a right triangle.

OpenStudy (shadowfiend):

So the next step is, do a quick sketch on paper of what the triangle would look like given that 8 and 10 are similar in length and 3 is very short. Does it look like it has an obtuse angle?

OpenStudy (anonymous):

no

OpenStudy (shadowfiend):

Right. So it would be an acute triangle.

OpenStudy (anonymous):

i kind of remember something that my teacher said about it that if its equal then its right if its less its acute if its more then obtuse is that right?

OpenStudy (shadowfiend):

That sounds right.

OpenStudy (anonymous):

ok thank you